OpenStudy (anonymous):

i am a parent who is lost only one question left and no clue here is the question ming rolls a number cube, tosses a coin and chooses a card from two cards marked a and b. if an even number and head appears ming wins no matter which card is chose. otherwise lashonda wins. find P(ming wins)

OpenStudy (tkhunny):

1) This problem cannot be solved very specifically, since we don't quite know the nature of the "number cube". 2) If we assume a standard die, it has 1, 2, 3, 4, 5, 6, that's 50% even and 50% odd. 3) Are you sure you presented the problem exactly? It seems like the cards are useless.

OpenStudy (anonymous):

please help me i am lost and dont know this stuff

OpenStudy (tkhunny):

p(Even) + p(Odd) = 0.5 + 0.5 = 1 p(Heads) + p(Tails) = 0.5 + 0.5 = 1 (p(Even) + p(Odd))*(p(Heads) + p(Tails)) = p(Even)p(Heads) + p(Even)p(Tails) + p(Odd)p(Heads) + p(Odd)p(Tails) Examine EACH piece: p(Even)p(Heads) = 0.5*0.5 = 0.25 ==> Ming Wins! p(Even)p(Tails) = 0.5*0.5 = 0.25 ==> Lashonda Wins! p(Odd)p(Heads) = 0.5*0.5 = 0.25 ==> Lashonda Wins! p(Odd)p(Tails) = 0.5*0.5 = 0.25 ==> Lashonda Wins! The way the problem is stated, that's all there is. It has nothing to dowith the Cards. If you would like to present a better problem statement, we can talk.